What is the quickest way to solve this? I know you can make a 16 punnett square but that takes long.Among the crosses below, which will give a 1:1 ratio of genotypes? A) AABB x aabb B) AaBb x AaBb C) AaBb x aabb D) AaBB x aaB
Answer #1

This problem can be solved by the probability method. First, calculate a probability for each monohybrid cross separately and then use the probability method to find the combined probability.

A) AA x aa = only one type of genotype is produced i.e. Aa

so the probability of Aa genotype = 4/4 = 1

BB x bb = also produces the same type of genotype i.e Bb

So the probability of the genotype AaBb = 1

B)Aa xAa genotype ratio is 1:2:1(AA:Aa:aa)

Probability of AA = 1/4, Probability of Aa= 2/4, probability of aa = 1/4

Bb xBb genotype ratio = 1:2:1 (BB:Bb:bb) probability of genotypes can be calculated same as Aa x Aa

The ratio of genotypes in both cases is 1:2:1 so it is not possible to get 1:1 ratio

AABB (1/4*1/4 = 1/16)

AaBB (2/4*1/4,= 2/16)

aaBB (1/4 * 1/4 = 1/16)

AABb (1/4*2/4 = 2/16)

AAbb (1/4*1/4 = 1/16)

AaBb (2/4*2/4 = 4/16)

aabb (1/4* 1/4 = 1/16)

aaBb (1/4*2/4 = 2/4)

Aabb(2/4* 1/4 =2)

genotype ratio is 1:2:1:2:1:4:1:2:2

C) Aa x aa = 2 Aa and 2 aa

Bb x bb = 2 Bb and 2bb

Possible genotypes = AaBb = 2/4 * 2/4 = 4/16

aaBb = 2/4 *2/4 = 4/16

AaBb = 2/4 * 2/4 = 4/16

aabb = 2/4 * 2/4 = 4/16

Genotype ratio = 4:4:4:4 = 1:1:1:1

D) Aa x aa =2 Aa and 2 aa

Probability of each genotype = 2/4

BB x BB = all BB probaility = 1(4/4)

Possible genotypes =

probability of AaBB = 2/4 * 4/4 = 8/16

probability of aaBB = 2/4 * 4/4 = 8/16

The ratio of the genotypes = 8:8 = 1:1

(E)AA x aa = all Aa

bb xBB = all Bb

So only one genotype is possible AaBb

C and D will give 1: 1 ratio, A and E give the only a single type of genotype.

You can also use these tricks for calculation

Crossa | Поточої Homozygous helerozygous x heleso- domilant pomozygous zygous. Homozygous. Hererorygous. I recessive Aa x Aa